OpenStudy (anonymous):

5/6 + n+4/3n-4 Simplify please the answer is supposed to be: 21n+4/6(3n-4) can you show me how to get there?

OpenStudy (anonymous):

\[\frac{n+4}{3 n-4}+\frac{5}{6}=\frac{6(n+4)+5(3 n-4)}{6(3 n-4)}=\frac{24+6 n+-20+15 n}{6(3 n-4)}\]\[=\frac{(21 n+4)}{6(3 n-4)}\]